ROTATIONAL ANALYSIS OF THE VIBRATIONAL GROUND STATE OF DIMETHYL ETHER, CH_{3}OCH_{3}

\underline{P. GRONER}, Department of Chemistry, University of Missouri - Kansas City, Kansas City, MO 64110; S. ALBERT, E. HERBST and F. C. DE LUCIA, Department of Physics, Ohio State University, Columbus, OH 43210.

An effective rotational Hamiltonian^{a} was used to analyze the rotational transitions in the vibrational ground state of dimethyl ether. Microwave^{b} and mm-wave^{c} measurements from the literature were combined with new measurements between 100 and 550 GHz in a global fit of all four torsional substates. Frequencies between 8 and 550 GHz were fit for transitions involving energy levels with J up to 40 and K_{a} up to 9. Only 22 spectroscopic parameters were necessary to fit 1499 frequencies to experimental precision (dimensionless standard deviation 0.67). The following parameters were determined in the least-squares fit: ρ = 0.21665(19), β = 8.426(27) deg., parameters equivalent to the rotational, quartic and sextic distortion constants, the internal energy tunneling parameters ϵ_{01} = −3.0392(43) MHz and ϵ_{02} = 0.0019(17) MHz and three tunneling constants related to the ”rotational” constants.

aP. Groner, submitted.

bF. J. Lovas, H. Lutz and H. Dreizler, J. Phys. Chem. Ref. Data (1979) 8, 1057-1107; J. R. Durig, Y. S. Li and P. Groner, J. Mol. Spectrosc. (1976) 62, 159-174.

cW. Neustock, A. Guarnieri, J. Demaison and G. Wlodarczak, Z. Naturforsch., Part A (1990) 45, 702-706.